### v1

# Analogical Structure, Minimal Contextual Cues and Contrastive Distractors: Input Design for Sample-Efficient Linguistic Rule Induction

Abstract
Large language models achieve strong performance through training on vast datasets. Can analogical paradigm organization enable lightweight models to match this performance with minimal data? We develop a computational approach implementing three cognitive-inspired principles: analogical structure, contrastive learning, and minimal contextual cues. We test this approach with structured completion tasks where models identify correct sentence completions from analogical patterns with contrastive alternatives. Training lightweight models (BERT+CNN, $0.5M$ parameters) on only one hundred structured examples of English causative/inchoative alternations achieves $F1=0.95$, outperforming zero-shot GPT-o3 ($F1=0.87$). Ablation studies confirm that analogical organization and contrastive structure improve performance, consistently surpassing randomly shuffled baselines across architectures. Cross-phenomenon validation using unspecified object alternations replicates these efficiency gains, confirming approach robustness. Our results show that analogical paradigm organization enables competitive linguistic rule learning with orders of magnitude less data than conventional approaches require.
Analogical Structure, Minimal Contextual Cues and Contrastive Distractors: Input Design for Sample-Efficient Linguistic Rule Induction
Chunyang Jiang1,2, Paola Merlo1,2 1Idiap Research Institute, Switzerland 2University of Geneva, Switzerland Firstname.Lastname@unige.ch
1 Introduction
Analogical reasoning, recognizing structural relationships across different contexts, enables efficient learning by abstracting patterns from minimal examples and transferring knowledge to novel situations (lake2017building). However, this principle has faced persistent scalability challenges across computational approaches: early symbolic systems required extensive knowledge engineering (forbus1995mac), neural word embeddings showed inconsistent performance (mikolov2013efficient), and even transformer-based models like BERT show mixed capabilities (ushio-etal-2021-bert; thrush2020investigating). While recent systems achieve strong performance (yasunaga2024large; jiayang-etal-2023-storyanalogy; wijesiriwardene-etal-2023-analogical), they still require substantial computational resources or show inconsistent results across complexity levels.
Rather than scaling analogical reasoning systems, we operationalise analogical principles through strategic input organization. Research shows that processing constraints can optimize attention allocation (christiansen2016now), while contrastive learning frameworks show how systematic positive-negative comparisons improve discriminative learning (chen2020simple; he2020momentum). This suggests that structural organization, not architectural scaling, may unlock analogical efficiency.
[…]
In this paper, we develop a computational approach that organizes input data into analogical paradigms through three cognitive-inspired strategies: (i) Analogical Structure, systematic paradigmatic relationships supporting pattern recognition; (ii) Contrastive Learning, systematically designed distractors enabling discriminative learning; and (iii) Minimal Contextual Cues, subtle annotations providing essential semantic information without explicit labeling.
We evaluate this setup using English verb alternations (Levin93) through structured completion tasks based on Blackbird Language Matrices (merlo-2023-blackbird). These tasks require identifying correct sentence completions from systematically organized contexts with contrastive alternatives. Figure 1 illustrates our approach: contexts encode analogical mappings across parallel paradigms (e.g., Man:Dice :: Exploring:Mat), requiring models to abstract structural and semantic role relationships rather than memorize surface forms.
Results demonstrate sample efficiency: lightweight models (BERT+CNN, $\sim 0.5M$ parameters) trained on only $100$ structured examples achieve $F1=0.95$, outperforming zero-shot GPT-o3 ($F1=0.87$). Cross-phenomenon validation confirms robustness beyond causative constructions.
Scope: We operationalise computational algorithms inspired by cognitive principles without claiming to model human cognition directly.
Our contributions are: (i) a computational approach that implements analogical paradigm organization for sample-efficient linguistic rule learning; (ii) empirical evidence that this approach enables competitive performance with minimal training examples; (iii) systematic demonstration that lightweight models can match zero-shot LLM performance on linguistic rule induction tasks.  [^1] Code, data, and replication steps will be made publicly available upon publication.
2 Data and Task Design
We operationalize three cognitive-inspired principles through systematic paradigm organization using Blackbird Language Matrices (merlo-2023-blackbird), structured completion tasks that test systematic linguistic rule learning. This section describes our task framework, linguistic domain, and distractor design methodology.
2.1 Task Framework
Our completion tasks organize linguistic data into analogical paradigm structures. Each instance contains a structured Context comprising two parallel paradigms encoding analogical relationships (Pattern A $\leftrightarrow$ Pattern B), supplemented with minimal contextual cues that provide semantic scaffolding without explicit labeling. It also contains a systematic Answer Set: One correct completion that maintains established patterns, and six systematically designed distractors that violate specific structural, semantic, or paradigmatic constraints.
Success requires integrating our three cognitive principles. Models must recognize analogical mappings across paradigms (analogical structure), discriminate between valid and invalid transformations (contrastive learning), and use semantic scaffolding from minimal soft annotations (minimal contextual cues).
Type Example P S R Tests
Correct The mat rolled into a pillow $\times$ $\times$ $\times$ All principles (multi-layered analogical reasoning)
$R$ The explorer rolled into a pillow $\times$ $\times$ $\checkmark$ Semantic roles (minimal contextual cues)
SC-$R$ The explorer was into a pillow $\times$ $\checkmark$ $\checkmark$ Syntactic structure (contrastive learning)
SCRS The mat rolled the explorer $\times$ $\checkmark$ $\checkmark$ Argument structure (syntax + semantics)
PC-$R$ The man rolled in the cup $\checkmark$ $\times$ $\checkmark$ Cross-paradigm consistency (analogical structure)
PSC-$R$ The man was in the cup $\checkmark$ $\checkmark$ $\checkmark$ Multiple constraints (analogical + contrastive)
PSC-RS The dice rolled the man $\checkmark$ $\checkmark$ $\checkmark$ Complex analogical mapping (all principles)
P=Paradigm, S=Structure, R=Role violations
Table 1: Systematic distractor taxonomy testing cognitive principles through hierarchical constraint violations. Each error type violates specific linguistic dimensions ($\checkmark$) to isolate analogical reasoning components.
2.2 Linguistic Test Domain: English Verb Alternations
Causative/Inchoative Alternations
[…]
This linguistic phenomenon provides an ideal test case because it requires analogical reasoning across different lexical items sharing the same structural pattern. It also needs to discriminate between valid and invalid argument structure transformations, and integrate semantic role information with syntactic patterns.
Cross-Phenomenon Validation
We validate the generalisability of the approach with bake-class verbs, a class of verbs exhibiting the unspecified object alternations (The chef baked a cake vs. The chef baked). In this class, the subject of both the transitive and the intransitive is an Agent. While the two classes share syntactic structure, they instantiate semantically distinct processes, testing whether our organizational principles generalize beyond specific semantic classes.
2.3 Analogical Paradigm Organization
Figure 1 illustrates our systematic organization approach. Each context follows a $2\times 4$ paradigmatic structure.
Multi-Layered Analogical Structure
Contexts encode analogical relationships at multiple levels: Man:Dice :: Explorer:Mat in terms of Agent-Theme relationships, with identical structural transformations applied across different lexical content. This tests whether models can abstract relational patterns beyond surface similarity.
Minimal Contextual Cues
We provide semantic scaffolding through subtle annotations within each paradigm. For example, action descriptors indicate agentivity (The player did it) while state descriptions suggest thematic roles (The ball was on the floor). These cues clarify semantic relationships without explicit grammatical labeling.
Contrastive Learning Through Systematic Distractors
Each answer set implements contrastive learning by providing systematic negative examples alongside each correct completion. Table 1 defines our hierarchical error taxonomy designed to test specific cognitive principles through constraint violations.
Each distractor violates exactly one constraint dimension while preserving others, enabling precise evaluation of rule component acquisition. Role errors (RR) test semantic understanding, structural errors (SC-RR, SCRS) test syntactic knowledge, and paradigm errors (PC-RR, PSC-RR, PSC-RS) test analogical consistency across contexts.
Notice that distractors are contextually inappropriate rather than inherently ungrammatical. For example, The explorer rolled into a pillow is grammatical but violates the analogical mapping established in the context.
2.4 Data Generation and Availability
Our systematic generation process uses template-based sentence creation with controlled lexical variation and seed alternation pairs, following established verb classifications (Levin93). [^2] Verb class and seed sentences are detailed in Appendix B
3 Experiments
We systematically evaluate our analogical paradigm organization approach through controlled experiments testing three core claims: (i) our organizational principles enable sample efficiency compared to unstructured baselines, (ii) individual components contribute systematically to performance, and (iii) lightweight models trained with structured data can match zero-shot reasoning-capable LLM performance on linguistic rule learning. The performance on BLM puzzle completion is used as a proxy for a model’s level of successful linguistic rule learning.
3.1 Experimental Design
Ablation Framework:
We test five systematic conditions that isolate organizational components while controlling information content.
Base is the complete paradigmatic organization with analogical structure and contextual cues. Shuffled has identical content in random order, (tests organizational vs. content effects). NoAnalogy removes first paradigm (tests analogical structure contribution). NoSoftCue removes contextual annotations (tests scaffolding contribution). Transposed preserves analogical patterns but changes spatial arrangement ( $C_{\text{Trans}}=C_{\text{Base}}^{\top}$).  [^3] See the formal definitions in Appendix A.1.
Data Configurations:
Type I uses same verbs across paradigms ($3000$ examples, 80:10:10 splits, direct analogical mapping). Type II uses different verbs across paradigms ($15000$ examples, same splits, tests abstraction). Cross-phenomenon uses unspecified object alternations with bake-class verbs (validates generalisability).
Scale Testing
Training sizes range from $10$ to $2700$ examples for lightweight models; and $300$-example test sets (same as for Type I test data, and $300$ random examples from Type II test set) for LLMs. This balances computational feasibility with statistical reliability for cross-model comparisons.
3.2 Models and Training
Lightweight Models
We use several models of representation. We embed each sentence using BERT444bert-base-multilingual-cased(devlin-etal-2019-bert) as primary model, RoBERTa555xlm-roberta-base (liu-etal-2019-roberta) and ELECTRA666electra-base-discriminator (clark-etal-2020-electra) for encoder comparison. We reason that using a weaker base clarifies contributions attributable to data organization rather than encoder capabilities.
As for architecture types, we use previously tested parameter-efficient neural models (CNN and FFNN in nastase-merlo-2023-grammatical for comparison).(1) CNN captures localized patterns and positional relationships; (2) FFNN processes concatenated embeddings through fully-connected layers, integrating distributed information across the entire context. Both architectures maintain minimal parameter counts ($\sim 0.5M$ parameters) compared to the encoder ($\sim 110M$), isolating the influence of data organization from architectural capacity. Both models output an answer embedding given the concatenated embeddings of the sentences from the context, evaluated against answer candidates using a max-margin objective with cosine similarity. Training: $120$ epochs, learning rate $0.001$, Adam Optimizer, batch size $100$, early stopping (patience $10$). All experiments use random seed $42$ with results averaged over $3$ independent runs. More model and training details are provided in the Appendix.
Figure 2: Zero-shot prompt.
Large Language Models
We evaluate eight advanced models including reasoning-capable systems. Reasoning models: deepseek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability), gpt-o3, gpt-o3-mini,qwq-32B; Standard models: deepseek-V3 (deepseekai2024deepseekv3technicalreport), llama-3.3-70B-Instruct, llama-3.2-3B-Instruct, qwen3-32B. The temperaturetemperature is $0.1$ where applicable (standard models), and the max_tokensmax\_tokens is $2046$.  [^7] Details on LLM configurations are in Appendix A.1.
Figure 3: F1 Performance as a function of training size. (A) Comparison of model architectures with BERT embeddings on Base structure organization. (B) Impact of data organization using the best architecture.
We test zero, one and five-shot prompting with (w-CoT) and without (wo-CoT) chain-of-thought reasoning (wei2021finetuned; kojima2022large). We use a puzzle-solving prompt that parallels our lightweight model setting. A zero-shot example of it is in Figure 2. The LLM first sees the context, generates a provisional answer, then compares it to the answer set before generating the best final sentence (instead of answer key label), and it also generates the hypotheses considered. Any final choice not in the answer set is flagged as a system error (ERR). This prompt mimics our learning objective using pretrained models and presents a more challenging task than standard multiple-choice questions. Randomizing the options in the answer set and asking to generate sentences instead of option labels help debias LLM’s token bias (e.g., A/B/C/D) in multiple-choice-format questions (zheng2024large). This uniform puzzle-solving format enables direct comparisons across model scales and training paradigms.
Evaluation
F1 scores balance precision and recall across 7 answer choices. Responses not matching answer set options are flagged as system errors for LLMs. Multi-factor ANOVA tests Structure ×× Model ×× Shots ×× CoT interactions ($p<0.05$).
3.3 Results
Figure 4: Isolated contributions of organizational components to best model performance. (A)  Impact of analogical organisation, comparing Base against NoAnalogy and Shuffled. (B)  Impact of implicit soft annotations, comparing Base with NoSoftCue and Transposed variants.
We present results from our best-performing architecture (CNN) across all experimental conditions with lightweight models. While both CNN and FFNN demonstrate similar qualitative trends, CNN consistently outperforms FFNN across training regimes, likely due to its capacity to capture local sequential patterns critical for analogical mapping.
3.3.1 Sample Efficiency Validation
Figure 3 confirms our core finding: lightweight models with Base organization achieve $F1=0.95$ with only $100$ examples, outperforming our best zero-shot reasoning model GPT-o3 ($F1=0.87$). Performance stabilizes at 100012001000-1200 examples across architectures. Panel B shows the impact of data organization, where Base structure outperforms Shuffled and Transposed arrangements. Identical content produces different learning trajectories, confirming that organizational structure, not information quantity, drives efficiency gains.
3.3.2 Component Analysis
Figure 5: Reasoning models (top row: deepseek-R1, gpt-o3, gpt-o3-mini, qwq-32B) vs. standard models (bottom row: llama-3.3-70B-Instruct, llama-3.2-3B-Instruct, qwen3-32B, deepseek-V3) across data structures. Red dotted line shows small model baseline (F1=0.98). Reasoning models consistently outperform standard models, but struggle to match structured lightweight model performance in zero-shot settings.
Figure 4 validates systematic contributions from individual organizational components. While all structures eventually converge to similar performance levels at larger training sizes, analogical structure (Base vs. NoAnalogy vs.Shuffled) shows slightly more consistent advantages in the 5001,500500-1,500 example range where data efficiency matters most. Contextual cues (Base vs. NoSoftCue vs. Transposed) provide additional but smaller benefits.
This subtle differences in learning curves suggest that while both components contribute to learning, modern neural architectures can compensate for deficiencies in either component as training sizes increases.
3.3.3 LLM Comparison
Figure 5 confirms our third contribution: lightweight models trained on $1000$ structured examples outperform all tested LLMs in zero-shot conditions, including state-of-the-art reasoning models. Multi-factor ANOVA reveals significant main effects for Structure, Model, and Shots, with minimal CoT impact [^8] Complete results on Multi-factor ANOVA appear in Appendix D.2 (Table 3). . This confirms this task engages pattern recognition and rule application rather than multi-step reasoning. [^9] Complete results on LLMs performance are in the Appendix D.3 (Table 4).
Reasoning vs. Standard Models
Reasoning models start with high zero-shot performance ($F1=0.8+$) but plateau quickly, while standard models begin lower (F1=0.10.5F1=0.1-0.5) but show steeper improvement curves, making them more sensitive to organizational structure.
Organizational Structure Patterns
Shuffled consistently performs worst across all models, confirming that random organization disrupts learning. Base and NoAnalogy emerge as most effective, often performing comparably. Notably, NoAnalogy sometimes outperforms Base in zero-shot settings for reasoning models, suggesting analogical structure becomes less crucial when models already possess strong reasoning capabilities. However, standard models benefit most from organizational effects that have greater room for improvement.
3.4 Error Analysis and Patterns
Figure 6: Error analysis across data organizations. Lightweight models (left) show robust performance with minimal errors ($3.3\%$) across all structures. LLMs (right, GPT-o3 zero-shot) show structure-dependent error patterns following our cognitive-inspired error taxonomy (Table 1). Shuffled structures particularly disrupt LLMs’ analogical reasoning, causing increased multi-constraint violations.
Figure 6 shows organizational structure affects error patterns across different model types. Lightweight models are robust ($errors<3.3\%$) across all organizational conditions, demonstrating consistent learning regardless of structural variations. LLMs show structure-dependent error patterns following our error taxonomy. Disrupted analogical structure (Shuffled condition) causes increased multi-constraint violations, with correct responses dropping from $55\%$ (Base) to $24\%$ (Shuffled) for gpt-o3. This shows that even reasoning-capable models rely heavily on structural organization for linguistic rule learning tasks.
3.5 Additional Analysis
Cross-Phenomenon Generalization
Figure 7: Cross-phenomenon validation using unspecified object alternations (bake-class verbs). Results show F1 performance using BERT+CNN architecture on Type-II data, averaged over 3 runs.
Figure 7 confirms Base >> Shuffled hierarchy replicates with bake-class verbs. This validates generalisability beyond causative constructions.
Figure 8: (A) Seen type vs. Unseen Type performance across training sizes.
Cross-Type Generalization
Figure 8 demonstrates successful rule abstraction: models achieve robust cross-type performance by 100200100-200 examples, though all conditions eventually converge to high performance with sufficient training data. This confirms learning of structural relationships rather than surface patterns.
Discussion
Our systematic evaluation validates all three core claims. The finding that lightweight structured models ($F1=0.98$) exceed even gpt-o3’s reasoning capabilities ($F1=0.87$) demonstrates that systematic input organization can achieve superior linguistic rule learning with reduced computational requirements. The consistent organizational effects across both reasoning and standard LLMs suggest that analogical paradigm organization addresses fundamental challenges in pattern recognition that persist even in the most advanced current systems. The contrasting error patterns between model types suggest that organizational benefits operate through different mechanisms, with accelerated learning for lightweight models and improved error discrimination for LLMs. The cross-phenomenon and cross-type evaluation confirms that our approach captures general principles rather than construction-specific optimizations.
These findings indicate that structured input data organization is complementary to performance gains for linguistic rule learning beyond architectural advances or reasoning capabilities.
4 Related Work
Structured Data for Sample Efficiency
Achieving competitive performance with reduced training needs remains a central challenge in NLP. Recent work emphasizes how reorganizing raw text into structured or hierarchically arranged contexts can significantly enhance model learning (liu2024enhancing; sourati2024arn). “Structured data” often refers to knowledge graphs, table data, databases (jiang-etal-2023-structgpt; li2023resdsql). liu2024enhancing instead reorganize text into hierarchical formats (e.g., Scope-Aspect-Description) to mimic human knowledge consolidation, thereby improving task performance in question-answering. Existing sample efficiency approaches in NLP include data selection (albalak2024a), data augmentation (kumar-etal-2020-syntax; dai-etal-2023-auggpt), curriculum learning (bengio2009curriculum; xu-etal-2020-curriculum), among others. However, they typically focus on scaling or sequencing existing data rather than fundamental reorganization principles. Our work instead targets linguistic rule learning through structure in a completion tasks inspired by cognitive principles.
Analogical Learning in Neural Language Models
Analogical reasoning is the ability to perceive and use similarities on relational patterns rather than surface features (gentner2012analogical). Neural models have shown strong capabilities of analogical reasoning to generalize rules, such as word-level analogies (mikolov2013efficient; brown2020language; ushio-etal-2021-bert; webb2023emergent), concept analogies through analogical structure abduction (yuan-etal-2023-beneath), narrative-based or story-level analogies (nagarajah2022understanding; jiayang-etal-2023-storyanalogy; sourati2024arn), or analogical prompting where LLMs generate relevant exemplars or knowledge before solving the problem (yasunaga2024large). These approaches, like ours, suggest that explicit structural cues improve analogical processing.
Contrastive Learning in NLP
Contrastive learning has emerged as a powerful principle for neural representation learning. Systematic comparison between positive and negative examples enhances discriminative learning (chen2020simple; he2020momentum). However, existing contrastive approaches typically require large-scale training data and focus on representation learning rather than rule acquisition. Our systematic distractor design similarly implements a form of contrastive learning by organizing examples that emphasize critical distinctions between alternations. This enables models to learn from the structured contrasts between valid and invalid linguistic transformations.
Linguistic Rule Learning in Neural Networks
Neural approaches to linguistic rule learning are challenging despite recent advances. Neural language models demonstrate emergent syntactic capabilities (linzen-etal-2016-assessing; gulordava-etal-2018-colorless; hewitt-manning-2019-structural; mueller-etal-2024-context) but requires large-scale training. kann-etal-2019-verb examine alternations in transformer embeddings with limited success, thrush2020investigating find that systematic selectional preferences require careful training design. LLMs also struggle on metalinguistic tasks (thrush-etal-2024-strange). merlo-2023-blackbird introduces Blackbird Language Matrices (BLM) as structured evaluation paradigms for systematic rule assessment, enabling controlled testing of linguistic competence. However, they show limited efficiency in data-lean settings. Our work addresses these efficiency challenges through analogical paradigm organization, building on the BLM framework to achieve competitive argument structure learning with minimal training examples while demonstrating robustness across alternation types.
5 Conclusion
We investigated whether computational approaches inspired by cognitive principles can enable sample-efficient linguistic rule learning. Our analogical paradigm organization approach operationalises three organizational strategies, analogical structure, contrastive learning, and minimal contextual cues, through systematic input structuring rather than architectural scaling.
Our results demonstrate substantial sample efficiency gains across multiple evaluation dimensions. Lightweight models (BERT+CNN, $\sim 0.5M$ parameters) trained on only $100$ structured examples achieved $F1=0.95$, outperforming zero-shot GPT-o3 ($F1=0.87$) and stabilizing performance with just 100012001000-1200 training instances. Component analysis confirmed that each organizational principle contributes systematically, with analogical structure showing the largest effect. Cross-phenomenon validation with bake-class verbs and cross-type generalization experiments confirmed that benefits extend beyond construction-specific optimizations.
These findings have practical and theoretical implications. Practically, our approach enables competitive linguistic rule learning with much fewer resources than conventional LLM approaches. Theoretically, we demonstrate that systematic input organization can achieve efficiency gains that complement rather than require architectural scaling. This suggests that cognitive-inspired data structuring is a distinct optimization dimension.
Future work should explore cross-linguistic validation, integration with pre-training objectives, and automated structure discovery methods. This research shows that computational approaches inspired by cognitive principles can provide practical advances in sample-efficient linguistic rule learning.
6 Limitations
While our approach demonstrates efficiency gains, it is not without limitations.
Expert Dependency
Our structured paradigms require expert linguistic knowledge to design appropriate analogical mappings and systematic distractors. Although we demonstrate cross-phenomenon generalisability with bake-class verbs, scaling this approach to diverse linguistic phenomena would require either linguistic expertise or automated methods for discovering optimal organizational structures.
Pre-training Attribution
Our lightweight models rely on pre-trained encoders (BERT, RoBERTa, ELECTRA) that already encode linguistic knowledge from unstructured data. This creates attribution challenges in isolating whether performance gains stem purely from our organizational principles or from interactions with pre-existing linguistic representations. While our Base vs. Shuffled comparisons control for this confound, the fundamental attribution question remains.
Language and Phenomenon Scope
Our evaluation was exclusively on English verb alternations, a specific subset of linguistic phenomena at the syntax-semantics interface. Cross-linguistic validation and extension to other grammatical constructions (e.g., morphological patterns, syntactic transformations) are needed to establish broader applicability of our organizational principles.
Computational Comparison Framework
Our comparison between trained lightweight models and zero-shot LLMs may not reflect equivalent computational investments. While we demonstrate superior sample efficiency, a fairer comparison might involve fine-tuning LLMs on equivalent structured data or comparing against few-shot learning with similar computational budgets.
Task Specificity
Our approach targets pattern completion tasks that may favour analogical reasoning. Extension to diverse NLP applications (e.g., generation, reasoning) would provide stronger evidence for the general usability of our organizational principles.
Ethics
Our research presents minimal ethical concerns. Our synthetically generated data focuses on grammatical verb alternations rather than semantic content, minimizing bias propagation risks from LLM training data. Code, data, and experimental procedures will be made publicly available upon publication.
Appendix A Experimental Setup Details
A.1 Ablation Design Formalization
We systematically manipulate context structures to isolate organizational components. Given a full context matrix $C_{\text{full}}\in\mathbb{R}^{m\times n}$ and transformation matrix $T\in{0,1}^{m\times n}$, transformed contexts are computed as:
$C_{\text{trans}}=C_{\text{full}}\odot T$
Transformation Matrices
  • NoAnalogy Context
    $T_{\text{af}}=\begin{bmatrix}0&0&0&0\\ 1&1&1&0\end{bmatrix}$
  • NoSoftCue Context
    $T_{\text{scf}}=\begin{bmatrix}1&0&0&1\\ 1&0&0&0\\ \end{bmatrix}$
  • For Transposed Context, it simply involves taking the transpose of the full context matrix.
    $C_{\text{trans}}=C_{\text{full}}^{T}$
A.2 Statistical Analysis Framework
Multi-factor ANOVA Design
Structure (5 levels) $\times$ Model (8 levels) $\times$ Shots (3 levels) $\times$ CoT (2 levels) $\times$ Data Type (2 levels)
Generalization Gap Metric
$\text{GenGap}(t,s)=\frac{1}{|M|}\sum_{m\in M}\left[F1(m,s,\text{seen},t)-F1(m,s,\text{unseen},t)\right]$ (1)
where $t$ = training size, $s$ = structure type, $M$ = model set, and smaller gaps indicate better generalization.
Appendix B Linguistic Phenomena
B.1 Causative/Inchoative Alternation (Roll-Class Verbs)
The class of roll verbs, as categorised by Levin93, comprises verbs expressing dynamic actions with inherent motion characteristics. This class shows systematic alternation between:
  • Intransitive: Theme as subject (e.g., ‘The ball rolled”)
  • Transitive: Agent causes Theme motion (e.g.,‘ The player rolled the ball”)
This syntactic-semantic interface represents a complex mapping challenge for computational models.
Verb Inventory
The verbs within this class include: bounce, coil, drift, drop, float, glide, move, revolve, roll, rotate, slide, spin, swing, turn, twirl, twist, whirl, and wind.
B.2 Cross-Phenomenon Validation (Bake-Class Verbs)
Unspecified object alternation with bake-class verbs provides syntactically similar but semantically distinct validation:
  • Transitive: ‘The chef baked a cake”
  • Intransitive:‘ The chef baked” (object understood)
This tests whether organizational benefits generalize across alternation types rather than being phenomenon-specific optimizations.
Appendix C Model Implementation
C.1 Encoder Comparison
Figure 9: F1 performance as a function of training size, comparing encoder architectures (BERT, RoBERTa, ELECTRA) with CNN architecture on Base structure. Train and test on type I data.
We employ BERT (bert-base-multilingual-cased) as our primary encoder to emphasize contributions attributable to data organization rather than encoder capabilities, as evidenced by Figure 9.
C.2 Architecture Specifications
Convolutional Neural Network (CNN)
  • Input: Array of embeddings, size $7\times 768$
  • Convolution layers: Three 2D convolutional layers
    • Kernel size: $3\times 3$, Stride: 1, No dilation
  • Output: Fully connected layer compressing to size $768$
  • Function: Localizes sequential patterns for analogical mapping
Feed-Forward Neural Network (FFNN)
  • Input: Concatenated sentence embeddings, size $7\times 768$
  • Architecture: Three fully connected layers
  • Compression: $7\times 768\rightarrow 3.5\times 768\rightarrow 3.5\times 768\rightarrow 768$
  • Function: Distributed pattern integration across entire context
C.3 Training Configuration
Models are trained over $120$ epochs with learning rate $0.001$ using the Adam optimizer and batch size $100$. Early stopping with patience $10$ prevents overfitting. All models employ a max-margin objective using cosine similarity:
$\mathcal{L}=\sum_{\mathbf{e}_{i}\in\mathcal{A}}\max\left(0,1+\cos(\mathbf{e}_{i},\mathbf{e}_{\text{pred}})-\cos(\mathbf{e}_{c},\mathbf{e}_{\text{pred}})\right)$ (2)
where $\mathbf{e}_{c}$ is the correct answer embedding, $\mathbf{e}_{i}$ are incorrect distractor embeddings, and $\mathbf{e}_{\text{pred}}$ is the predicted completion. Models are run across three independent trials with seed $42$ to ensure statistical reliability.
C.4 Computational Environment
Hardware Configuration
  • Workstation: HP PAIR Workstation Z4 G4 MIT
  • Processor: Intel Xeon W-2255
  • RAM: 64 GB
  • GPU: MSI GeForce RTX 3090 VENTUS 3X OC, 24 GB GDDR6X
Software Environment
  • Ubuntu: 22.04, Python: 3.11.5, CUDA: 11.8.0
  • PyTorch: 2.2.2, Transformers: 4.39.1, scikit-learn: 1.5.0
  • NumPy: 1.25.0, pandas: 1.4.3
Appendix D Complementary Experimental Results
D.1 Large Language Model Specifications
We evaluate eight advanced models across four families with varying parameter scales. This systematic comparison allows assessment of how organizational structure affects different model capacities and architectural approaches.
Model Family Variant Parameters Context Type Release
Reasoning Models
DeepSeek deepseek-R1-0528 670B 64K Reasoning 2025-05
OpenAI gpt-o3 Undisclosed 128K Reasoning 2025-04
OpenAI gpt-o3-mini Undisclosed 128K Reasoning 2025-01
Alibaba qwq-32B 32B 32K Reasoning 2025-03
Standard Models
DeepSeek deepseek-V3-0324 671B 64K Standard 2025-03
Meta llama-3.3-70B-Instruct 70B 128K Standard 2024-12
Meta llama-3.2-3B-Instruct 3B 128K Standard 2024-09
Alibaba qwen3-32B 32B 128K Standard 2025-04
Table 2: Large language models evaluated in experiments. Context window indicates maximum input sequence length in tokens.
D.2 Statistical Analysis Results
Complete multi-factor ANOVA results demonstrating significant main effects for Structure, Model, and Shots across all experimental conditions.
Factors F p
C(Structure) 9284.541 $p<0.05$
C(Model) 9581.462 $p<0.05$
C(Num_Shots) 14125.142 $p<0.05$
C(CoT) 9.562 $p<0.05$
C(Data_Type) 191.786 $p<0.05$
C(Structure):C(Model) 313.487 $p<0.05$
C(Structure):C(Num_Shots) 170.116 $p<0.05$
C(Model):C(Num_Shots) 664.883 $p<0.05$
C(Structure):C(CoT) 2.125 0.0601
C(Model):C(CoT) 3.625 $p<0.05$
C(Num_Shots):C(CoT) 5.399 $p<0.05$
C(Structure):C(Data_Type) 115.123 $p<0.05$
C(Model):C(Data_Type) 37.116 $p<0.05$
C(Num_Shots):C(Data_Type) 34.928 $p<0.05$
C(CoT):C(Data_Type) 0.078 0.7795
C(Structure):C(Model):C(Num_Shots) 55.286 $p<0.05$
C(Structure):C(Model):C(CoT) 1.644 $p<0.05$
C(Structure):C(Num_Shots):C(CoT) 2.265 $p<0.05$
C(Model):C(Num_Shots):C(CoT) 1.889 $p<0.05$
C(Structure):C(Model):C(Data_Type) 13.123 $p<0.05$
C(Structure):C(Num_Shots):C(Data_Type) 17.556 $p<0.05$
C(Model):C(Num_Shots):C(Data_Type) 11.177 $p<0.05$
C(Structure):C(CoT):C(Data_Type) 0.368 0.8709
C(Model):C(CoT):C(Data_Type) 0.214 0.9824
C(Num_Shots):C(CoT):C(Data_Type) 6.551 $p<0.05$
C(Structure):C(Model):C(Num_Shots):C(CoT) 0.899 0.7079
C(Structure):C(Model):C(Num_Shots):C(Data_Type) 7.100 $p<0.05$
C(Structure):C(Model):C(CoT):C(Data_Type) 0.919 0.6045
C(Structure):C(Num_Shots):C(CoT):C(Data_Type) 1.748 0.0657
C(Model):C(Num_Shots):C(CoT):C(Data_Type) 1.221 0.2531
C(Structure):C(Model):C(Num_Shots):C(CoT):C(Da… 0.944 0.6086
Table 3: Multi-ANOVA tests on data (Structure, Data_Type) and experimental (Model, Num_Shots, CoT) factors.
D.3 Comprehensive Performance Results
Complete results across all models, structures, data types, and shot configurations with and without Chain-of-Thought reasoning.
Model Config Base Noanalogy Nosoftcue Shuffled Transpose
deepseek-R1 (0-shot, w-CoT) 0.743 ± 0.048 0.692 ± 0.037 0.711 ± 0.049 0.614 ± 0.043 0.655 ± 0.050
deepseek-R1 (0-shot, wo-CoT) 0.719 ± 0.063 0.707 ± 0.018 0.720 ± 0.020 0.641 ± 0.030 0.681 ± 0.035
deepseek-R1 (1-shot, w-CoT) 0.857 ± 0.004 0.876 ± 0.048 0.802 ± 0.017 0.788 ± 0.012 0.851 ± 0.028
deepseek-R1 (1-shot, wo-CoT) 0.848 ± 0.009 0.884 ± 0.054 0.802 ± 0.005 0.765 ± 0.021 0.873 ± 0.071
deepseek-R1 (5-shot, w-CoT) 0.884 ± 0.052 0.994 ± 0.004 0.875 ± 0.072 0.871 ± 0.073 0.927 ± 0.076
deepseek-R1 (5-shot, wo-CoT) 0.883 ± 0.058 0.911 ± 0.065 0.866 ± 0.069 0.871 ± 0.073 0.859 ± 0.007
deepseek-V3 (0-shot, w-CoT) 0.537 ± 0.193 0.859 ± 0.089 0.397 ± 0.163 0.445 ± 0.107 0.196 ± 0.130
deepseek-V3 (0-shot, wo-CoT) 0.487 ± 0.409 0.912 ± 0.102 0.630 ± 0.320 0.483 ± 0.410 0.139 ± 0.177
deepseek-V3 (1-shot, w-CoT) 0.729 ± 0.015 0.961 ± 0.008 0.634 ± 0.019 0.620 ± 0.056 0.544 ± 0.082
deepseek-V3 (1-shot, wo-CoT) 0.725 ± 0.011 0.963 ± 0.008 0.632 ± 0.019 0.604 ± 0.033 0.509 ± 0.092
deepseek-V3 (5-shot, w-CoT) 0.895 ± 0.063 0.994 ± 0.002 0.853 ± 0.058 0.703 ± 0.021 0.828 ± 0.100
deepseek-V3 (5-shot, wo-CoT) 0.963 ± 0.007 0.995 ± 0.003 0.899 ± 0.013 0.716 ± 0.036 0.867 ± 0.076
gpt-o3 (0-shot, w-CoT) 0.852 ± 0.008 0.754 ± 0.031 0.844 ± 0.053 0.688 ± 0.041 0.838 ± 0.012
gpt-o3 (0-shot, wo-CoT) 0.884 ± 0.057 0.754 ± 0.022 0.840 ± 0.064 0.700 ± 0.030 0.834 ± 0.015
gpt-o3 (1-shot, w-CoT) 0.891 ± 0.087 0.995 ± 0.012 0.858 ± 0.076 0.783 ± 0.053 0.869 ± 0.066
gpt-o3 (1-shot, wo-CoT) 0.911 ± 0.082 0.990 ± 0.008 0.834 ± 0.031 0.785 ± 0.046 0.894 ± 0.079
gpt-o3 (5-shot, w-CoT) 0.939 ± 0.072 0.998 ± 0.004 0.884 ± 0.094 0.858 ± 0.075 0.922 ± 0.086
gpt-o3 (5-shot, wo-CoT) 0.925 ± 0.078 1.000 ± 0.000 0.842 ± 0.019 0.830 ± 0.033 0.938 ± 0.085
gpt-o3-mini (0-shot, w-CoT) 0.754 ± 0.026 0.629 ± 0.028 0.606 ± 0.044 0.431 ± 0.041 0.641 ± 0.034
gpt-o3-mini (0-shot, wo-CoT) 0.745 ± 0.047 0.644 ± 0.070 0.636 ± 0.052 0.410 ± 0.014 0.573 ± 0.101
gpt-o3-mini (1-shot, w-CoT) 0.809 ± 0.015 0.863 ± 0.055 0.673 ± 0.024 0.542 ± 0.060 0.701 ± 0.063
gpt-o3-mini (1-shot, wo-CoT) 0.812 ± 0.023 0.866 ± 0.059 0.647 ± 0.068 0.551 ± 0.044 0.730 ± 0.068
gpt-o3-mini (5-shot, w-CoT) 0.894 ± 0.076 0.989 ± 0.013 0.774 ± 0.042 0.638 ± 0.044 0.767 ± 0.053
gpt-o3-mini (5-shot, wo-CoT) 0.881 ± 0.086 0.995 ± 0.005 0.764 ± 0.045 0.622 ± 0.053 0.786 ± 0.047
llama-3.2-3B-Instruct (0-shot, w-CoT) 0.093 ± 0.035 0.478 ± 0.033 0.109 ± 0.014 0.118 ± 0.015 0.061 ± 0.033
llama-3.2-3B-Instruct (0-shot, wo-CoT) 0.105 ± 0.042 0.501 ± 0.034 0.112 ± 0.018 0.113 ± 0.015 0.066 ± 0.037
llama-3.2-3B-Instruct (1-shot, w-CoT) 0.524 ± 0.079 0.761 ± 0.031 0.388 ± 0.044 0.279 ± 0.053 0.283 ± 0.049
llama-3.2-3B-Instruct (1-shot, wo-CoT) 0.532 ± 0.055 0.757 ± 0.032 0.374 ± 0.039 0.271 ± 0.057 0.269 ± 0.036
llama-3.2-3B-Instruct (5-shot, w-CoT) 0.456 ± 0.071 0.882 ± 0.045 0.330 ± 0.051 0.317 ± 0.043 0.308 ± 0.031
llama-3.2-3B-Instruct (5-shot, wo-CoT) 0.491 ± 0.063 0.890 ± 0.038 0.313 ± 0.041 0.340 ± 0.051 0.319 ± 0.026
llama-3.3-70B-Instruct (0-shot, w-CoT) 0.304 ± 0.083 0.700 ± 0.015 0.340 ± 0.034 0.337 ± 0.026 0.238 ± 0.110
llama-3.3-70B-Instruct (0-shot, wo-CoT) 0.298 ± 0.087 0.687 ± 0.017 0.309 ± 0.021 0.318 ± 0.018 0.223 ± 0.106
llama-3.3-70B-Instruct (1-shot, w-CoT) 0.774 ± 0.021 0.845 ± 0.046 0.515 ± 0.036 0.491 ± 0.042 0.592 ± 0.073
llama-3.3-70B-Instruct (1-shot, wo-CoT) 0.783 ± 0.047 0.783 ± 0.080 0.499 ± 0.015 0.447 ± 0.046 0.594 ± 0.066
llama-3.3-70B-Instruct (5-shot, w-CoT) 0.964 ± 0.055 0.996 ± 0.001 0.875 ± 0.084 0.737 ± 0.035 0.952 ± 0.045
llama-3.3-70B-Instruct (5-shot, wo-CoT) 0.942 ± 0.071 0.976 ± 0.051 0.870 ± 0.091 0.734 ± 0.021 0.931 ± 0.056
qwen3-32B (0-shot, w-CoT) 0.786 ± 0.122 0.596 ± 0.222 0.736 ± 0.029 0.441 ± 0.009 0.538 ± 0.074
qwen3-32B (0-shot, wo-CoT) 0.659 ± 0.088 0.662 ± 0.059 0.752 ± 0.069 0.429 ± 0.016 0.535 ± 0.065
qwen3-32B (1-shot, w-CoT) 0.808 ± 0.011 0.797 ± 0.018 0.792 ± 0.018 0.532 ± 0.014 0.729 ± 0.046
qwen3-32B (1-shot, wo-CoT) 0.811 ± 0.021 0.788 ± 0.009 0.800 ± 0.012 0.525 ± 0.028 0.740 ± 0.046
qwen3-32B (5-shot, w-CoT) 0.839 ± 0.003 0.896 ± 0.070 0.830 ± 0.017 0.636 ± 0.049 0.815 ± 0.007
qwen3-32B (5-shot, wo-CoT) 0.833 ± 0.009 0.917 ± 0.067 0.842 ± 0.063 0.619 ± 0.021 0.808 ± 0.010
qwq-32B (0-shot, w-CoT) 0.776 ± 0.035 0.724 ± 0.023 0.747 ± 0.026 0.504 ± 0.018 0.672 ± 0.076
qwq-32B (0-shot, wo-CoT) 0.752 ± 0.060 0.732 ± 0.011 0.731 ± 0.036 0.462 ± 0.042 0.652 ± 0.076
qwq-32B (1-shot, w-CoT) 0.843 ± 0.010 0.821 ± 0.015 0.793 ± 0.007 0.598 ± 0.042 0.762 ± 0.056
qwq-32B (1-shot, wo-CoT) 0.843 ± 0.009 0.822 ± 0.015 0.792 ± 0.011 0.600 ± 0.047 0.755 ± 0.048
qwq-32B (5-shot, w-CoT) 0.863 ± 0.006 0.863 ± 0.004 0.831 ± 0.017 0.691 ± 0.021 0.846 ± 0.021
qwq-32B (5-shot, wo-CoT) 0.864 ± 0.002 0.863 ± 0.003 0.829 ± 0.009 0.680 ± 0.017 0.836 ± 0.027
Table 4: Performance (F1 scores in % $\pm$ standard deviation) of all LLMs validated across all configurations.
Appendix E Additional Analyses
E.1 Learning Dynamics Analysis
Figure 10: Reduction in error rates as training size increases for the Base structure (train and test on type I data). Dashed lines represent smoothed trends using LOWESS smoothing.
Figure 10 tracks error reduction as training progresses with the Base structure. ParadigmChange-RoleReplace errors show the most dramatic decrease, suggesting paradigm-level understanding develops earlier than fine-grained role distinctions. The Base structure facilitates more efficient error reduction across all types, providing better implicit feedback about linguistic rules consistent with error-driven learning theories.
E.2 Generalization Analysis
Figure 11: Generalization gap analysis for base and shuffled structures. Smaller gaps indicate better generalization. Base structure facilitates better transfer of linguistic rules to new contexts across all training sizes. Smoothing uses centered rolling window (window size=3) with confidence bands representing standard error propagation.
Figure 11 demonstrates that Base structure maintains consistent generalization patterns while Shuffled shows erratic performance in mid-range training sizes. At small training sizes, both structures show negative gaps, suggesting models avoid overfitting to surface lexical patterns. Base structure shows more stable and efficient generalization throughout the learning process.
E.3 Cross-Phenomenon Validation
Cross-phenomenon validation with bake-class verbs confirms that organizational benefits (Base >> Shuffled) replicate across different linguistic constructions, validating generalisability beyond causative alternations and demonstrating that our approach captures general organizational principles rather than construction-specific optimizations.